OpenStudy (anonymous):

PB has endpoints at –2 and 12. What is the coordinate of its midpoint?

OpenStudy (mertsj):

Add the endpoints and divide by 2

OpenStudy (anonymous):

The midpoint of CD is E(–1, 0). One endpoint is C(5, 2). What are the coordinates of the other endpoint?

OpenStudy (mertsj):

\[(\frac{x _{1}+x _{2}}{2},\frac{y _{1}+y_{2}}{2})\]

OpenStudy (mertsj):

\[\frac{5+x _{2}}{2}=-1\] \[\frac{2+y _{2}}{2}=0\]

OpenStudy (anonymous):

(-7,-2)

OpenStudy (anonymous):

-7 , -2 is the answer to what @Mertsj is this person wrong

OpenStudy (mertsj):

(-7,-2) is correct.

OpenStudy (anonymous):

to the second question

OpenStudy (mertsj):

The midpoint of CD is E(–1, 0). One endpoint is C(5, 2). What are the coordinates of the other endpoint? The answer to this question is (-7,-2)